use the given confidence interval limits to find the point estimate and the margin of error for the equation: 0.772 is less than p and p is less than 0.776
Answer by stanbon(75887)

use the given confidence interval limits to find the point estimate and the margin of error for the equation: 0.772 is less than p and p is less than 0.776
-----------------------
The point estimate is the average of the confidence interval limits:
----
p-hat = (0.772+0.776)/2 = 0.774
------
Margin of Error = 0.776-0.774 = 0.002
===============
Cheers,
Stan H.
===============